7,15,31,63, ...
The first term in the sequence above is 7, and each term after the first is
Determined by multiplying the preceding term by m and then adding p. What is the value of m?
(A) 1
(B) 2
(C) 3
(D) 4
(E) 9

Answer by richwmiller(17219)

m=2 and p=1
2*7+1=15
2*15+1=31
2*31+1=63